OpenStudy (anonymous):

HELP ME !! Garrett has deposited $519 in a savings account that earns interest at a rate of 4.2% compounded daily. What will the account balance be in 17 years? $540.80 $1,044.52 $1,059.83 $2,179.0

jimthompson5910 (jim_thompson5910):

FV = PV*(1+r/n)^(n*t) FV = 519*(1+0.042/365)^(365*17) FV = 1059.82895074734 FV = 1059.83